OpenStudy (anonymous):

Hi I know it's late but I was wondering if someone could help me real quickly with a problem I'm having . Using the points in the corner creat the following equation. 1. Determine the equation of the line,in standard form,that will get your spacecraft from the launch area to point A Launch area x-coordinate 1 , y-coordinate 2 Point A x-coordinate 2 ,y-coordinate 4 I know the slope is M=2

OpenStudy (p0sitr0n):

general form of a line: y=mx+b y=2x+b as you found to find b just replace a pair of (x,y) into the equation 4=2*2+b b turns out to be 0 so its y=2x

OpenStudy (p0sitr0n):



OpenStudy (anonymous):

Thank so much I have been struggling on this question for most of the week and the teachers won't help saying that their nothing they can really do .